### v1

# Q*: Improving Multi-step Reasoning for LLMs with Deliberative Planning

Abstract
Large Language Models (LLMs) have demonstrated impressive capability in many nature language tasks. However, the auto-regressive generation process makes LLMs prone to produce errors, hallucinations and inconsistent statements when performing multi-step reasoning. In this paper, we aim to alleviate the pathology by introducing Q*, a general, versatile and agile framework for guiding LLMs decoding process with deliberative planning. By learning a plug-and-play Q-value model as heuristic function, our Q* can effectively guide LLMs to select the most promising next step without fine-tuning LLMs for each task, which avoids the significant computational overhead and potential risk of performance degeneration on other tasks. Extensive experiments on GSM8K, MATH and MBPP confirm the superiority of our method.
1 Introduction
Large Language Models (LLMs) have exhibited impressive capabilities in solving various reasoning tasks encoded in nature languages, including math word problems [1, 2, 3, 4, 5, 6], code generation [7, 8, 9] and planning [10, 11, 12]. Unfortunately, even the most advanced LLMs still face significant challenges and are prone to introduce errors, hallucinations and inconsistent statements as the number of reasoning steps grows due to their auto-regressive nature [13, 14]. In fact, the auto-regressive generation process of LLMs can be characterized by “System 1" [15], a mode of thought which is fast, instinctive but less accurate. Most of recent works focus on improving LLMs’ System 1 capability by (1) constructing sophisticated prompts with extensive expertise to trigger the potential capacities of LLMs without modifying their parameters [16, 17, 18, 19], (2) fine-tuning LLMs with massive task-specific corpus at the price of significant computational burdens and the potential risk of performance degeneration on other tasks [5, 6, 20, 21], or (3) training reward models to rank the candidate responses [22, 23, 4, 24].
On the other hand, solving complex reasoning problems requires more in-depth, deliberative and logical thinking steps, i.e., the “System 2" mode [15]. Taking solving math word problems as an example, any incorrect intermediate reasoning step (e.g., calculation errors, mis-interpretations) can potentially lead to incorrect final answers. Prior attempts [25, 26, 27, 28] for enhancing System 2 reasoning capability includes performing deliberation with basic tree search algorithms (e.g., BFS or DFS), Monte Carlo Tree Search (MCTS) [29], and A* [30]. Nonetheless, the utility functions used in these methods often require laborious expertise to design for each specific task, which are difficult to be extended to new scenarios. Furthermore, deliberation with MCTS would require significant number of rollouts before finding high-quality responses when solving the problems with many reasoning steps, which substantially slows down the overall decoding process.
In light of this, we propose Q*, a general, versatile and agile framework for improving the multi-step reasoning capability of LLMs with deliberative planning. Different from the existing deliberation methods, our method does not rely on domain knowledge to design the heuristic function. Besides, by leveraging plug-and-play Q-value models as heuristic function, our Q* can effectively solve various tasks via guiding LLMs to select the most promising next step without fine-tuning LLMs beforehand, which avoids the significant computational overhead and potential risk of performance degeneration in other tasks. Finally, Q* considers only a single step when performing deliberation, which is much cheaper than complete rollouts in MCTS. Specifically, the main contributions of our work are summarized as follows:
  • We formalize the multi-step reasoning of LLMs as a Markov Decision Process (MDP) where the state is the input prompt and the reasoning steps generated so far, the action is the next step of reasoning and and the reward measures how well the task is solved.
  • We present several general approaches to estimate the optimal Q-value of state-action pairs, i.e., offline reinforcement learning, best sequence from rollout and completion with stronger LLMs. It is noteworthy that our methods only need the ground truth of training problems and can be easily applied to various reasoning tasks without modification.
  • We cast solving multi-step reasoning tasks as a heuristic search problem, where the objective is to find the most proper reasoning trace with maximum utility. Built upon A* search, our deliberation framework, Q*, leverages plug-and-play Q-value models as heuristic function and guides LLMs to select the most promising next reasoning step in best-first fashion.
  • We conduct extensive experiments on math reasoning and code generation tasks, demonstrating that Q* can significantly improve the multi-step reasoning capability of LLMs.
2 Related Works
LLM alignment.
Alignment is an important technique to prevent the output of LLMs deviates from human’s expectation. Supervised Fine-Tuning (SFT) is probably the most fundamental alignment approach that directly minimizes the cross-entropy loss between the output and ground-truth. Reinforcement learning from Human Feedback (RLHF) [31], on other hand, firstly learns a reward model (RM) from human preferences and then optimizes the SFT model with reinforcement learning algorithms to maximize the cumulative rewards from RM. Direct Preference Optimization (DPO) [32] aligns LLMs directly according to the ranking information from human feedback without explicitly learning RM. Recently, Aligner [33] came out as a model-agnostic alignment method by learning to re-write LLMs’ output. Compared to these methods, our Q* achieves alignment with distinct merits. Different from SFT and Aligner, Q* does not rely on massive human generated corpus which are expensive to collect; different from RLHF and DPO, our Q* does not modify the parameters of LLM, which avoid potential performance loss on other tasks.
Enhancing LLMs with planning.
Tree-of-thoughts (ToT) [25] improves the LLMs’ reasoning capability by exploring the intermediate steps towards problem solving with basic tree-search algorithms. In the same vein, A* search and MCTS are applied to serve as a planning technique to enhance the performance of LLMs when solving challenging reasoning problems [26, 27, 28, 34]. Unfortunately, the utility function used in these methods is either constructed from LLMs’ feedback (e.g., [25, 27]), which could be highly-inaccurate in complex problems, or specific to each individual task (e.g., [28, 34]). Moreover, planning with MCTS often requires to perform costly rollout, which can significantly slow down the overall decoding process. In contrast, our Q* relies on ground-truth solely to train value model and can be easily applied to various reasoning tasks without modification. Besides, we consider only a single step each time in Q*, which is much cheaper than complete rollout in MCTS-based methods.
LLMs for math reasoning & code generation.
Math reasoning and code generation require LLMs to perform multi-step reasoning on relations, quantities and logics which are inherently challenging. Current techniques include 1) prompt engineering which triggers the potential capacities of LLMs with sophisticated prompts [16, 17, 18, 19, 35, 36]. However, constructing such prompt needs extensive expertise and case-by-case tuning, which is difficult to generalize to different tasks; 2) Fine-tuning LLMs with massive math/code corpus [5, 6, 20, 21, 8, 9, 37], which usually at the price of significant computational burden and may compromise the performance on other tasks; 3) training RMs/verifiers to rank the candidate solutions without providing any guidance in intermediate steps [22, 23, 4, 24]. Differently, our Q* leverages a plug-and-play Q-value model to direct the deliberation process of LLMs, which effectively provides guidance for each intermediate step without modifying the parameters of LLMs. Moreover, by casting multi-step reasoning of LLMs as a heuristic search problem, our method can be generalized to various reasoning tasks without laborious prompt engineering.
3 Preliminary
3.1 Formulate the Multi-step Reasoning of LLMs as an MDP
Taking the question $q$ as input, the answer generation process of LLMs can be broken down into multiple reasoning steps, where the final answer sequence $\mathbf{a}$ can be treated as the concatenation of these $T$ single-step reasoning steps, formulated as $\mathbf{a}=[a_{1};a_{2};\dots;a_{T}]$. Each step can be a single line or fixed number of tokens outputted by LLMs. Under this perspective, we can conceptualize the multi-step reasoning process of LLMs as a Markov Decision Process (MDP) $\langle\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R},\gamma\rangle$, where the state $s_{t}\in\mathcal{S}$ denotes the partial reasoning trace already generated by timestep $t-1$ plus the input question (i.e., $s_{t}=[q;a_{1};\dots;a_{t-1}]$) with the special case $s_{1}=q$, the action $a_{t}\in\mathcal{A}$ denotes the next reasoning step generated by LLMs taking the current state $s_{t}$ as input, the deterministic state transition $\mathcal{T}$ from the current state $s_{t}$ to the next state $s_{t+1}$ is accomplished through a simple operation of concatenation, $\mathcal{R}$ is the reward function to measure how well the question is solved and $\gamma$ is the discount factor. The reward function is often outcome-based, i.e., it gives reward by comparing the final results with ground-truth. Finally, the policy $\pi_{\theta}$ is embodied by an LLM, which produces reasoning sequence conditioned on the input question:
$\pi_{\theta}(a_{t}|s_{t})=\text{LLM}(a_{t}|s_{t}),~{}\pi_{\theta}(\mathbf{a}|q)=\prod_{t=1}^{T}\pi_{\theta}(a_{t}|s_{t}).$ (1)
Given the MDP and LLM policy $\pi_{\theta}$, the value of state-action pair $(s_{t},a_{t})$ is given by a Q-function $Q^{\pi_{\theta}}(s_{t},a_{t})=\mathbb{E}_{\pi_{\theta}}\left[\sum_{t^{\prime}=t}^{T}\gamma^{T-t^{\prime}}\mathcal{R}(s_{t^{\prime}},a_{t^{\prime}})\right]$. The Q-function of an optimal policy $\pi^{*}$ is called optimal Q-function and satisfies the Bellman optimality equation:
$Q^{*}(s_{t},a_{t})=\mathcal{R}(s_{t},a_{t})+\gamma\max_{a_{t+1}\in\mathcal{A}}Q^{*}(s_{t+1},a_{t+1}),$ (2)
which gives the value of starting state $s_{t}$, taking action $a_{t}$ and then following the optimal policy $\pi^{*}$.
3.2 A* Search
A* [30] is an important heuristic search algorithm in deliberative planning [38], multi-agent pathfinding [39], and constraint reasoning [40]. Originally, A* was proposed for finding the shortest path from source $s$ to goal $g$ in path planning problems. It associates each frontier vertex $n$ with a value $f(n)=g(n)+h(n)$, where $g(n)$ is the accumulated cost from source $s$ and $h(n)$ is a heuristic value that estimates the cost of the shortest path from $n$ to goal $g$. The algorithm adopts a best-first search strategy, i.e., in each iteration it always picks the vertex with minimum $f$-value to explore until reaching the goal. When the heuristic $h(\cdot)$ is admissible, A* guarantees to find the optimal path.
Figure 1: Overview of Q*. (a): the deliberation process of Q*. Each state is associated with a $f$-value which is the weighted sum of the aggregated utility (cf. Eq. (4)) and the heuristic value (cf. Eq. (5)). (b-d): estimating optimal Q-value with fitted-Q-iteration, rollout and completion with stronger LLMs.
4 Q*: A General, Versatile and Agile Deliberation Framework for LLMs
Most of modern LLMs generate natural languages in an auto-regressive way, i.e., predict the next token in a sequence given the previously generated tokens (cf. Eq. (1)). Therefore, when applied to multi-step reasoning problem, LLMs can potentially introduce errors, hallucinations and inconsistent statements in subsequent reasoning trace if any previous step is incorrect, which may fail to solve the problem. Indeed, given the fact that LLMs produce each token with fixed computation, there is no way to investigate more computational power to solve difficult problems. Simply put, LLMs cannot perform in-depth deliberation which is essential for solving complex reasoning problem.
We address the issue by presenting Q*, a general, versatile and agile deliberation framework based on A* to effectively guide LLMs to select the most promising next step when perform multi-step reasoning without costly fine-tuning LLMs for each task beforehand. In more detail, we cast finding the most proper reasoning sequence for a given problem as a heuristic search process, where each state $s_{t}$ is associated with a $f$-value estimating how much utility will be attained if we expand $s_{t}$:
$f(s_{t})=g(s_{t})+\lambda h(s_{t}),$ (3)
where $g(s_{t})$ is the aggregated utility from the initial state $s_{1}$, $h(s_{t})$ is the heuristic value measuring the quality of the final result derived from $s_{t}$ and $\lambda$ is a coefficient to balance $g(s_{t})$ and $h(s_{t})$.
Specifically, we propose to use process-based reward function $\mathcal{R}_{P}$ that encodes the prior knowledge or preference of the reasoning task to compute the aggregated utility $g(s_{t})$. That is,
$g(s_{t})=\text{Agg}(\mathcal{R}_{P}(s_{1}),\dots,\mathcal{R}_{P}(s_{i}),\dots,\mathcal{R}_{P}(s_{t})),$ (4)
[…]
$f(s_{t})=g(s_{t})+\lambda\max_{a_{t}\in\mathcal{A}}Q^{*}(s_{t},a_{t}).$ (5)
Since maximizing over all possible next reasoning steps is intractable, in practice we restrict the alternatives to the top-K steps returned by LLM, i.e., $f(s_{t})=g(s_{t})+\lambda\max_{a_{t}\in\text{top-K}(\pi_{\theta}(\cdot|s_{t}))}Q^{*}(s_{t},a_{t})$.
4.1 Estimation of Optimal Q-value
A critical challenge of implementing Q* is to estimate the optimal Q-value of state-action pairs (cf. Eq.(5)) with a frozen LLM policy $\pi_{\theta}$ which could be suboptimal on the given reasoning problems. Therefore, we aim to learn a proxy Q-value model $\hat{Q}$ to approximate $Q^{*}$ from a dataset $D=\{q_{i},\{\mathbf{a}_{i_{j}}\}_{j=1}^{M}\}_{i=1}^{N}$, where $q_{i}$ is a training problem and $\mathbf{a}_{i_{j}}\sim\pi_{\theta}(\cdot|q_{i})$ is a trajectory sampled from the LLM policy $\pi_{\theta}$. Formally:
$\hat{Q}=\arg\min_{Q}\frac{1}{NMT}\sum_{i=1}^{N}\sum_{j=1}^{M}\sum_{a_{t}\in\mathbf{a}_{i_{j}}}\left(Q(s_{t},a_{t})-\hat{y}(s_{t},a_{t})\right)^{2},$ (6)
where $s_{t}=[q_{i};a_{1};\dots;a_{t-1}]$ is the partial reasoning trace by timestep $t-1$ in $\mathbf{a}_{i_{j}}$ and $\hat{y}(s_{t},a_{t})$ is the label that approximates the true optimal Q-value $Q^{*}(s_{t},a_{t})$.
In more detail, we effectively construct Q-value labels $\hat{y}(s_{t},a_{t})$ for question $q_{i}$ in the following ways:
Offline reinforcement learning.
Given the offline dataset $D$, we learn the proxy Q-value model $\hat{Q}$ using Fitted Q-iteration [42]. Specifically, for each iteration $\ell$, we construct Q-value label as:
$\hat{y}_{\ell}(s_{t},a_{t})=\begin{cases}\mathcal{R}(s_{t},a_{t})&t=T\\ \mathcal{R}(s_{t},a_{t})+\gamma\max_{a_{t+1}\in\text{top-K}(\pi_{\theta}(\cdot|s_{t+1}))}\hat{Q}_{\ell-1}(s_{t+1},a_{t+1})&\text{otherwise}\end{cases},$ (7)
where $\hat{Q}_{\ell-1}$ is the proxy Q-value model learned in iteration $\ell-1$. After that, we train a new proxy model $\hat{Q}_{\ell}$ according to Eq. (6). Such two phases will be alternated for $L$ iterations, and we use $\hat{Q}_{L}$ as the proxy Q-value model when performing deliberation.
Learning from rollout.
Staring from state-action pair $(s_{t},a_{t})$, we first perform random rollout or MCTS [27, 26] with policy $\pi_{\theta}$ for enough iterations, forming a trajectory pool $P$. After that, we use the best reasoning sequence with the highest accumulated reward to construct Q-value label:
$\hat{y}(s_{t},a_{t})=\mathcal{R}(s_{t},a_{t})+\max_{\begin{subarray}{c}\tau\sim P\\ (s_{t^{\prime}},a_{t^{\prime}})\in\tau\end{subarray}}\left[\sum_{t^{\prime}=t+1}^{T}\gamma^{T-t^{\prime}}\mathcal{R}(s_{t^{\prime}},a_{t^{\prime}})\right].$ (8)
Approximating optimal policy with stronger LLMs.
By exploiting the fact that the optimal Q-values are derived from an optimal policy, one can estimate the optimal Q-value of state-action pair $(s_{t},a_{t})$ by completing the trajectory with another stronger LLM $\pi_{\theta^{*}}$ (e.g., GPT-4) if available:
$\hat{y}(s_{t},a_{t})=\mathcal{R}(s_{t},a_{t})+\sum_{t^{\prime}=t+1}^{T}\gamma^{T-t^{\prime}}\mathcal{R}(s^{*}_{t^{\prime}},a^{*}_{t^{\prime}}),$ (9)
where $s^{*}_{t^{\prime}}=[s_{t};a_{t};a^{*}_{t+1};\dots;a^{*}_{t^{\prime}-1}]$ and $a^{*}_{t^{\prime}}\sim\pi_{\theta^{*}}(\cdot|s^{*}_{t^{\prime}})$.
4.2 Deliberative Planning with A*
Once obtaining proxy Q-value model $\hat{Q}$, we can plug it to Eq. (5) to compute the $f$-value of each state and perform best-first search with A*. Alg. 1 illustrates the deliberative planning process.
Algorithm 1 Deliberative planning for LLMs with A*
1:question $q$, LLM policy $\pi_{\theta}$, proxy Q-value model $\hat{Q}$
2:$pq\leftarrow\{q\}$, $visited\leftarrow\emptyset$
3:while $pq\neq\emptyset$ do
4:     $s\leftarrow\arg\max_{s^{\prime}\in pq}f(s^{\prime})$
5:     $pq\leftarrow pq\backslash\{s\}$, $visited\leftarrow visited\cup\{s\}$
6:     if $s$ is a terminal state then return extract-answer($s$)
7:     for each $a\in\text{top-K}(\pi_{\theta}(\cdot|s))$ do
8:         $s^{\prime}\leftarrow[s;a]$
9:         if $s^{\prime}\notin visited$ then $pq\leftarrow pq\cup\{s^{\prime}\}$
Specifically, we maintain an open list $pq$ which initially contains the input question, and a closed list $visited$ to record the visited states. Each step we pick the state $s$ with the maximum $f$-value from the open list and expand it by querying the top-K alternatives with the LLM policy $\pi_{\theta}$. After that, the open list is updated and the process repeats until the terminal state is reached. Finally, we extract the answer part of the terminal state as the result.
5 Experiments
5.1 Experimental Settings
Datasets. We evaluate the effectiveness of Q* on two math reasoning and one code generation tasks, where the dataset statistics have been summarized in Table 1. 1) GSM8K [2] is a dataset of grade school math problems, where the solution is given in a one-line-per-step format with an exact numerical answer in the last line; 2) MATH [3] is a dataset consisting of math problems of high school math competitions, where the solutions are given in a format that mixes latex code and natural language; 3) MBPP [43] is an entry-level Python programming dataset, where the questions are coding challenges along with a test case that defines the function format. The solutions are Python code that is excepted to pass the pre-collected test cases of each question.
Table 1: Statistics of datasets.
Dataset GSM8K MATH MBPP
Domain Math reasoning Math reasoning Code generation
Training 5000 8000 374
Testing 1319 5000 500
Average Steps 4.5 11.0 7.0
Implementation Details. The implementation of Q* method mainly include three steps: 1) Q-value estimation. As discussed in Section 4.1, we propose several ways for estimating Q values and will illustrate the implementation details in the next paragraph; 2) Aggregated utility for each step (cf. Eq. (4)). For GSM8K dataset, we adopt a process reward model $\mathcal{R}_{P}$ trained on PRM800K [22] to provide intermediate signal for each reasoning step, and use $\min$ as the aggregation function; For MATH dataset, we set $g(s)=0$ for all state $s$ for fairness because PRM800K contains samples constructed from MATH testing set and the trained PRM could potentially have seen the correct answer of testing problems; For MBPP dataset, we tokenize the code generated so far with function tokenize.generate_tokens and give a penalty of -0.5 if TokenError is raised, which is often the case that there are mismatched delimiters (e.g., parentheses, quotation marks) and invalid indention in the code. We use $[-1]$ as the aggregation function to cancel the previous penalties since the code is generated on-the-fly and mismatched delimiters may be fixed in subsequent steps. 3) A* planning. For GSM8K and MATH datasets, we treat a single line outputted by the LLM as an action, while the action in MBPP is defined as a code snippet with 24 tokens when planning. Besides, we set $\lambda=1$ when computing $f$-values and expand a state with $K=6$ actions. Finally, following the common practise of Best-of-$N$, in all experiments we perform planning to collect $N=6$ trajectories for each question, and select the one with the maximum $f$-value as the final result for evaluation.
[…]
5.2 Experimental Results
GSM8K. For the comparison on GSM8K dataset, we select Llama-2-7b [44] as our base model, whose accuracy can achieve 65.2% after finetuning on MetaMath [5]. Then, we treat Llama-2-7b finetuned on MetaMath as policy $\pi_{\theta}$, and perform rollout to collect Q-value labels for training Q-value model (QVM). For aggregated utility, we train a process reward model (PRM) on PRM800K [22] to provide intermediate signal for each reasoning step. With a PRM/QVM in hand, traditional methods tend to treat either of them as a verifier the select the Best-of-$N$ trajectory or utilize them to perform PPO training of RLHF. As the results shown in Table 2, we can find that with the same PRM/QVM, using it for verification performs significantly better than using it for alignment. Further, in the comparison of planning-based methods, we can find that with the same QVM, Q* method with constant aggregated utility can still outperform Best-of-$N$ method. With the PRM trained on PRM800K determining whether the intermediate reasoning steps are correct, Q* method that combines PRM and QVM achieves the best performance among the methods based on the Llama-2-7b model, surpassing the performance of close-sourced ChatGPT-turbo [45].
MATH. As the results shown in Table 3, considering the weak performance of Llama-2-7b fine-tuned on MetaMath for the MATH dataset, we seek for two other stronger base models to evaluate the effectiveness of the Q* method. The one is Llama-2-7b fine-tuned on Synthetic Data [46], which is constructed following the instruction of scaling up the SFT data, and achieves 41.9% accuracy on MATH dataset, approaching the performance of GPT-4 [47]. Another base model is DeepSeek-Math-7b [48], which could be the most powerful open-source 7b model for math reasoning, achieving 50.8% accuracy in our evaluation. From the results shown in the second and third blocks of Table 3, we found that the Q* method can lead to further performance improvements compared to the Best-of-$N$ method. Additionally, the performance of Q* based on DeepSeek-Math-7b has already surpassed a series of closed-source models on the leaderboard of MATH dataset  [^1] https://paperswithcode.com/sota/math-word-problem-solving-on-math, such as Gemini Ultra (4-shot).
Table 2: Comparison of Q* and other baselines on GSM8K dataset.
Base Model SFT Alignment Verification Accuracy
GPT-3.5 (5-shot) [49] Unknown PPO (RM) [31] - 57.1%
ChatGPT-instruct (0-shot) [45] Unknown PPO (RM) [31] - 71.3%
ChatGPT-turbo (0-shot) [45] Unknown PPO (RM) [31] - 77.7%
GPT-4 (0-shot) [45] Unknown PPO (RM) [31] - 91.9%
GPT-4 (5-shot) [49] Unknown PPO (RM) [31] - 92.0%
Llama-2-7b (0-shot) - - - 49.5%
Llama-2-7b (0-shot) MetaMath[5] - - 65.2%
Llama-2-7b (0-shot) MetaMath[5] PPO (PRM) [31] - 67.2%
Llama-2-7b (0-shot) MetaMath[5] PPO (QVM) [31] - 67.6%
Llama-2-7b (0-shot) MetaMath[5] - Best-of-$N$ (PRM) [22] 72.1%
Llama-2-7b (0-shot) MetaMath[5] - Best-of-$N$ (QVM) [22] 74.5%
Llama-2-7b (0-shot) MetaMath[5] - Q* (QVM) 78.8%
Llama-2-7b (0-shot) MetaMath[5] - Q* (PRM+QVM) 80.8%
Table 3: Comparison of Q* and other baselines on MATH dataset.
Base Model SFT Alignment Verification Accuracy
GPT-3.5 (0-shot) [47] Unknown PPO (RM) [31] - 23.5%
GPT-4 (0-shot) [47] Unknown PPO (RM) [31] - 42.5%
Gemini Ultra (4-shot) [50] Unknown PPO (RM) [31] - 53.2%
Llama-2-7b (0-shot) - - - 7.9%
Llama-2-7b (0-shot) MetaMath[5] - - 20.0%
Llama-2-7b (0-shot) Synthetic Data[46] - - 41.9%
Llama-2-7b (0-shot) Synthetic Data[46] PPO (QVM) [31] - 42.5%
Llama-2-7b (0-shot) Synthetic Data[46] - Best-of-$N$ (QVM) [22] 46.8%
Llama-2-7b (0-shot) Synthetic Data[46] - Q* (QVM) 49.1%
DeepSeek-Math-7b (0-shot) Unknown PPO (QVM) [31] - 50.8%
DeepSeek-Math-7b (0-shot) Unknown PPO (QVM) [31] Best-of-$N$ (QVM) [22] 54.3%
DeepSeek-Math-7b (0-shot) Unknown PPO (QVM) [31] Q* (QVM) [22] 55.4%
MBPP. As for the comparison on MBPP dataset, we also choose the most powerful open-source LLM in the aspect of code generation, specifically CodeQwen1.5-7b-Chat, as our base model for evaluating the effectiveness of Q*. Following a similar procedure of math reasoning, we train a QVM for Q-value estimation and manually construct heuristic function as described in the previous part of implementation details. From the results shown in Table 4, we can find that Q* can still outperform Best-of-$N$ method in the aspect of code generation, and achieve 77.0% in MBPP dataset, which is still a promising performance in the leaderboard of MPBB  [^2] https://paperswithcode.com/sota/code-generation-on-mbpp.
Table 4: Comparison of Q* and other baselines on MBPP dataset.
Base Model SFT Alignment Verification Accuracy
GPT-3.5 Turbo (self-debug) [51] Unknown PPO (RM) [31] - 72.8%
GPT-4 (self-debug) [51] Unknown PPO (RM) [31] - 80.2%
CodeQwen1.5-7b-Chat (0-shot) Unknown PPO (QVM) [31] - 74.6%
CodeQwen1.5-7b-Chat (0-shot) Unknown PPO (QVM) [31] Best-of-$N$ (QVM) [22] 75.0%
CodeQwen1.5-7b-Chat (0-shot) Unknown PPO (QVM) [31] Q* (PRM+QVM) [22] 77.0%
6 Conclusion
Solving challenging multi-step reasoning problems requires LLMs to perform in-depth deliberation beyond auto-regressive token generation. In this paper, we present Q*, a general, versatile and agile deliberation framework for LLMs. Unlike existing deliberation methods which need extensive expertise to design a utility function for each specific task, Q* relies on ground-truth solely to train value model and can be easily applied to various reasoning tasks without modification. Moreover, by leveraging plug-and-play Q-value models as heuristic function, Q* can effectively guide LLMs to solve various tasks without fine-tuning LLMs beforehand, which avoid potential performance degeneration on other tasks. Finally, our Q* is agile because we consider only a single step each time rather than complete rollouts (e.g., simulation in MCTS). Extensive empirical evaluations on math reasoning and code generation tasks confirm the superiority of our method.